OpenStudy (anonymous):

SAT question: If x and t are positive numbers that satisfy the equation below, what is the value of x/t?

OpenStudy (anonymous):

\[\sqrt{x ^{2}-t ^{2}}=2t-x\] The answer is apparently 5/4. I would like an explanation.

OpenStudy (anonymous):

one sec!

OpenStudy (anonymous):

Ok! This one's fun.

OpenStudy (anonymous):

Think right off the bat with problems like this: "I NEED to get each variable on the same side" - so I want the t's on one side and the x's on another

OpenStudy (anonymous):

1) Kill the square root: \[\sqrt{x^2 - t^2}^2 = (2t-x)^2\]

OpenStudy (anonymous):

\[x^2 -t^2 = (2t-x)^2\]

OpenStudy (anonymous):

give it a stab from there

OpenStudy (anonymous):

Yes, but when I do that I always end up getting \[5t ^{2}=0\] because \[x ^{2}-t ^{2}=4t ^{2}+x ^{2}\]

OpenStudy (anonymous):

oh... wait. There's another term, duh

OpenStudy (anonymous):

ahh! yup! You gotta FOIL the right side

OpenStudy (anonymous):

so you get x^2 - t^2 = 4t^2 - 4tx + x^2

OpenStudy (anonymous):

Oh.. lol. Okay. So then ... 4x = 5t. then 4x/5t = 1, and x/t = 5/4 because you have to multiply by the reciprocal and both sides. Am I right?

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

*facepalm* Let's hope I don't do this on the exam

OpenStudy (anonymous):

good job. I hate missing stuff like that...happens ALL the time! :)

OpenStudy (anonymous):

Thanks!